# DialSim: A Real-Time Simulator for Evaluating Long-Term Multi-Party Dialogue Understanding of Conversational Agents

### v4

@@ Abstract
Recent advancements in Large Language Models (LLMs) have significantly enhanced the capabilities of conversational agents, making them applicable to various fields (e.g., education). Despite their progress, the evaluation of the agents often overlooks the complexities of real-world conversations, such as real-time interactions, multi-party dialogues, and extended contextual dependencies. To bridge this gap, we introduce DialSim, a real-time dialogue simulator. In this simulator, an agent is assigned the role of a character from popular TV shows, requiring it to respond to spontaneous questions using past dialogue information and to distinguish between known and unknown information. Key features of DialSim include assessing the agent’s ability to respond within a reasonable time limit, handling long-term multi-party dialogues, and evaluating performance under randomized questioning with LongDialQA, a novel, high-quality question-answering dataset. Our experiments using DialSim reveal the strengths and weaknesses of the latest conversational agents, offering valuable insights for future advancements in conversational AI. DialSim is available at https://dialsim.github.io/.
@@
Multi-Party Dialogue Understanding of Conversational Agents
@@ 1 Introduction
Figure 1: An overall process of DialSim. Gray speech bubbles indicate predetermined utterances from the script, and white speech bubbles indicate spontaneous questions asked during the simulation. Colored speech bubbles indicate the agent’s responses to the questions. (Left) An unanswerable question. (Center) A question that references a specific time. (Right) A multi-hop question that requires understanding past sessions (i.e., the Left and Center boxes). The question is asked in the format chosen by the user, either in a multiple-choice format or as an open-ended question.
@@
Recent advancements in Natural Language Generation (NLG) within Large Language Models (LLMs) have significantly enhanced the capabilities of conversational agents. These agents are now integral to various fields, including entertainment (Zhou et al., 2023; Chen et al., 2024) and education (Ait Baha et al., 2023; Waisberg et al., 2024), providing personalized interactions that cater to individual preferences and interests. As they continue to evolve and become more widely adopted, it is crucial to rigorously assess their performance in real-world scenarios to ensure they meet user expectations and function effectively.
@@
Traditionally, the evaluation of conversational agents has relied on qualitative assessments of their responses. This process typically involves human evaluators or LLMs judging the quality of an agent’s utterances (Adiwardana et al., 2020; Zhang et al., 2020; Roller et al., 2021; Shuster et al., 2022; Lee et al., 2023; Kim et al., 2024) or comparing responses between different agents on platforms like Chatbot Arena (Chiang et al., 2024). While these methods provide valuable insights into aspects such as naturalness and alignment with user instructions, they do not fully capture the complexities of real-world interactions.
@@
In practice, conversational agents face a variety of challenges: engaging in real-time interactions, managing multi-party conversations, and recalling information from past dialogues. These scenarios demand more comprehensive evaluation methods—ones that test an agent’s ability to respond within a reasonable time constraint, understand multi-party dialogue contexts, and reason across extended interactions. To meet this demand, we introduce DialSim, a real-time dialogue simulator designed to evaluate the long-term multi-party dialogue understanding of conversational agents.
@@
DialSim places the agent in the role of a main character within a TV show, engaging in extensive conversations based on the show’s scripted content (see Figure 1). During each session, a randomly selected character asks a randomly sampled question at an unpredictable time. The agent is evaluated on its ability to respond appropriately, relying solely on the dialogue history and acknowledging when it lacks sufficient information. This approach enables rigorous testing of dialogue comprehension in unpredictable, realistic scenarios. Additionally, the agent’s real-time interaction capabilities are assessed through time constraints for responses (e.g., 1s, 3s, 5s). To the best of our knowledge, this is the first work evaluating conversational agents under time constraints, introducing a novel dimension to agent performance assessment.
@@
In order to run DialSim, a dialogue script and corresponding question-answer pairs are required. For this purpose, we created LongDialQA, a new question-answering dataset derived from long-term multi-party dialogues. It comprises dialogues from popular TV shows (i.e., Friends, The Big Bang Theory, and The Office), spanning approximately 1,300 sessions over five years, totaling around 350,000 tokens. Each session includes more than 1,000 questions curated through two approaches: refining questions from a fan quiz website and generating complex questions using extracted temporal knowledge graphs. ChatGPT-4 (OpenAI, 2023a) assisted in refining questions and extracting knowledge graphs, with all outputs meticulously reviewed to ensure quality.
@@
LongDialQA also incorporates adversarial testing to rigorously challenge agents’ reliance on dialogue history rather than pre-trained knowledge. Since LLM-based agents may possess prior knowledge about the TV shows (see Appendix A), we developed adversarial tests that modify character names in two specific ways: by swapping their names with each other (e.g., Joey $\leftrightarrow$ Monica) or by assigning new names to them (e.g., Joey $\rightarrow$ John). These adversarial scenarios help verify that the agent’s responses are grounded in the contextual dialogue history rather than pre-trained knowledge.
Using DialSim, we evaluated the latest conversational agents, uncovering both their strengths and limitations. Our findings provide valuable insights for advancing conversational AI, emphasizing the need for robust, real-world evaluation frameworks.
@@ 2 Related Works
Conversational Agents Evaluation    Early evaluation methods for conversational agents often relied on reference-based metrics (e.g., BLEU Papineni et al. (2002), ROUGE Lin (2004), METEOR Banerjee and Lavie (2005)), which compare model outputs to gold dialogue references but often show weak correlation with human judgment Liu et al. (2016). In contrast, human evaluation—where human annotators assess coherence, factual correctness, consistency, and engagingness of the generated responses—provides reliable assessments Adiwardana et al. (2020); Zhang et al. (2020); Roller et al. (2021); Shuster et al. (2022); Lee et al. (2023), but it is costly and time-consuming.
@@
With the advent of LLMs, new evaluation approaches have emerged. These include having LLMs evaluate utterances directly Li et al. (2023); Kim et al. (2024) or employing platforms (e.g., Chatbot Arena Chiang et al. (2024)) where humans rank responses from different agents. Despite these advances, existing methods are still limited to qualitative assessments of utterances and fail to capture real-world conversational scenarios (e.g., real-time interaction, and long-term multi-party dialogue). To address these limitations, we propose a dialogue simulator, DialSim, designed to evaluate a conversational agent’s comprehensive dialogue understanding capabilities.
Long-Term Dialogue Datasets    A representative dataset for long-term dialogue is Multi Session Chat (Xu et al., 2022), which features up to five sessions per dialogue. This dataset, created through crowdsourcing, ensures high-quality dialogues; however, generating longer dialogues via crowdsourcing has remained challenging. To address this issue, Conversation Chronicles (Jang et al., 2023) was developed by leveraging an LLM to create longer and more comprehensive conversational datasets. More recently, LoCoMo (Maharana et al., 2024) was created using both LLMs and crowdsourcing; it evaluates dialogue comprehension of an agent through various tasks (e.g., event summarization) in long-term dialogues. In contrast to other datasets generated through crowdsourcing or LLMs, LongDialQA leverages TV show scripts, naturally providing extended, multi-party dialogues that evolve over time. Building on these unique features, DialSim simulates realistic, long-term interactions to evaluate agents.
@@
Datasets Based on the TV Show Scripts  While both TV show scripts and other dialogue datasets effectively capture dialogue characteristics, scripts offer a significant advantage due to their abundance and accessibility. This makes them particularly valuable for various dialogue understanding tasks such as question answering (QA) (Yang and Choi, 2019; Sang et al., 2022), coreference resolution (Chen and Choi, 2016; Chen et al., 2017; Zhou and Choi, 2018), relation extraction (Rashid and Blanco, 2018; Yu et al., 2020), and summarization (Gorinski and Lapata, 2015; Papalampidi et al., 2020; Chen et al., 2022). Notable datasets derived from scripts include FriendsQA (Yang and Choi, 2019) and TVShowGuess (Sang et al., 2022). FriendsQA treats each TV show scene as an independent conversation, with questions aiming to locate specific answer spans. TVShowGuess is a multiple-choice dataset requiring the identification of anonymized speakers in a scene based on prior context from earlier scenes. While many studies have utilized TV show scripts to create such datasets, only LongDialQA includes unanswerable questions and fully utilizes the extended context of scripts.
@@ 3.1.1 Script Preprocessing
The script we used includes 5 consecutive seasons per TV show, with each season containing approximately 20 episodes. Each episode is composed of multiple scenes (i.e., session). Each script includes not only utterances but also descriptions of characters’ actions and scenes, as well as metadata unrelated to the plot (e.g., names of writers and directors). We manually filtered out all irrelevant parts to create $Script_{pre}$, which contains only the conversations between characters. Additionally, since some of our questions involve time conditions (e.g., “Which friend wasn’t allowed to drive Monica’s Porsche in October 1994?”), we manually assigned a date to each scene in $Script_{pre}$ to provide time information to the agent. These dates were determined based on the contents of the conversations and the air dates of the episodes. The specific rules for date assignments are detailed in Appendix B. We then selected scenes involving the main character (i.e., Friends: Ross, The Big Bang Theory: Sheldon, The Office: Michael [^2] The characters with the most lines in each script were selected.) from $Script_{pre}$ and sequentially numbered them as sessions $S_{i}$. This process resulted in the final dialogue $\mathcal{D}=\{\mathcal{S}_{1},\mathcal{S}_{2},...,\mathcal{S}_{N}\}$.
Figure 2: The overall process of question generation based on fan quizzes. First, we crawled fan quizzes from the web (1). Then, we applied filtering and revision processes to the crawled data (2-a, b). From this, we created secondary versions of the questions by adding dates to each (3-a). Then, we mapped each question to the scenes by determining whether it is answerable in that scene or not (3-b). Finally, we applied character style transfer to make the questions more natural (4).
@@ 3.1.2 Fan Quiz-Based Question Generation
We utilized a fan quiz website FunTrivia [^3] https://www.funtrivia.com/ to generate our questions. Fan quizzes cover a range of difficulty levels and focus on major events from each episode, making them promising for evaluating dialogue comprehension. Figure 2 illustrates our process for generating questions using fan quizzes. We began by extracting episode-specific quizzes from the site. Since these quizzes were created by dedicated fans, many required knowledge unrelated to the dialogue itself (e.g., “What is the name of the actor who played the clerk?”). To filter out these questions, we first selected quizzes that could be answered by referencing $Script_{pre}$ using ChatGPT-4 (OpenAI, 2023a). [^4] Fan quizzes exist for each episode, so we annotated them based on $Script_{pre}$ and then matched them to the sessions of $\mathcal{D}$. Questions about scenes without the main character are unanswerable, enabling us to design rigorous tests. Additionally, ChatGPT-4 annotated the scenes that served as evidence for each question. These annotations were verified by the authors to ensure accurate filtering and scene-mapping.
@@
Figure 3: The overall process of question generation based on the temporal knowledge graph. We first extracted quadruples and constructed a temporal knowledge graph (1). Then, we generated questions based on this and mapped each question to the sessions by determining whether it was answerable in that session or not, similar to fan quiz-based questions (2-1, 2-2). Character style transfer was performed afterwards (3).
@@ 3.1.3 Temporal Knowledge Graph-Based Question Generation
Fan quizzes are useful for generating our questions, but since they are episode-specific and user-generated, the questions don’t span multiple episodes and their numbers are limited ($\sim$1K). To address this, we constructed a knowledge graph for each session and used it to generate questions. Initially, we used ChatGPT-4 to extract triples (i.e., [head, relation, tail]) from each session $S_{i}$ in $\mathcal{D}$. These triples were then refined by the authors. We employed 32 relations (e.g., girlfriend) derived from DialogRE (Yu et al., 2020), a high-quality dataset where human annotators manually extracted relations from Friends scripts, classifying relationships between characters into 37 categories. We adapted and modified these relations for our purpose. More details about the relations are provided in Appendix D.1. Finally, we combined the triples from each session with their respective dates to create a temporal knowledge graph (TKG) composed of quadruples (i.e., [head, relation, tail, date]).
@@ 3.1.4 Final Data Processing
Answer Choices Generation    To create multiple-choice questions, we carefully crafted a set of answer choices for each question. First, for all questions, we included a choice “(E) I don’t know.”, which agents must choose if the questions are unanswerable. For questions sourced from fan quizzes, the four answer choices were taken from the original quiz. The correct answers for these questions were the same as the original quiz, while the unanswerable questions were fixed to (E).
@@
Question Style Transfer    In LongDialQA, questions are rephrased to reflect each character’s unique tone, creating the impression that the characters themselves are asking the questions (e.g., Generic style: “How did Rachel buy her new boots?$\rightarrow$ Style of Joey Tribbiani from Friends: “Hey, how did Rachel manage to snag those killer boots, huh?”). This transformation is powered by ChatGPT-4, and subsamples are reviewed by the authors to ensure that the original intent was preserved. More examples of style-transferred questions for each character are in Appendix E.

### v5

# DialSim: A Real-Time Simulator for Evaluating Long-Term Multi-Party Dialogue Understanding of Conversation Systems

@@ Abstract
Recent advancements in Large Language Models (LLMs) have significantly enhanced the capabilities of conversation systems, making them applicable to various fields (e.g., education). Despite their progress, the evaluation of the systems often overlooks the complexities of real-world conversations, such as real-time interactions, multi-party dialogues, and extended contextual dependencies. To bridge this gap, we introduce DialSim, a real-time dialogue simulator. In this simulator, a conversation system is assigned the role of a character from popular TV shows, requiring it to respond to spontaneous questions using past dialogue information and to distinguish between known and unknown information. Key features of DialSim include assessing the system’s ability to respond within a reasonable time limit, handling long-term multi-party dialogues, and evaluating performance under randomized questioning with LongDialQA, a novel, high-quality question-answering dataset. Our experiments using DialSim reveal the strengths and weaknesses of the latest conversation systems, offering valuable insights for future advancements in conversational AI. DialSim is available at https://dialsim.github.io/.
@@
Multi-Party Dialogue Understanding of Conversation Systems
@@ 1 Introduction
Figure 1: An overall process of DialSim. Gray speech bubbles indicate predetermined utterances from the script, and white speech bubbles indicate spontaneous questions asked during the simulation. Colored speech bubbles indicate the system’s responses to the questions. (Left) An unanswerable question. (Center) A question that references a specific time. (Right) A multi-hop question that requires understanding past sessions (i.e., the Left and Center boxes). The question is asked in the format chosen by the user, either in a multiple-choice format or as an open-ended question.
@@
Recent advancements in Natural Language Generation (NLG) within Large Language Models (LLMs) have significantly enhanced the capabilities of conversation systems. These systems are now integral to various fields, including entertainment (Zhou et al., 2023; Chen et al., 2024) and education (Ait Baha et al., 2023; Waisberg et al., 2024), providing personalized interactions that cater to individual preferences and interests. As they continue to evolve and become more widely adopted, it is crucial to rigorously assess their performance in real-world scenarios to ensure they meet user expectations and function effectively.
@@
Traditionally, the evaluation of conversation systems has relied on qualitative assessments of their responses. This process typically involves human evaluators or LLMs judging the quality of the system’s responses (Adiwardana et al., 2020; Zhang et al., 2020; Roller et al., 2021; Shuster et al., 2022; Lee et al., 2023; Kim et al., 2024) or comparing responses between different systems on platforms like Chatbot Arena (Chiang et al., 2024). While these methods provide valuable insights into aspects such as naturalness and alignment with user instructions, they do not fully capture the complexities of real-world interactions.
@@
In practical scenarios, conversation systems must handle real-time interactions, manage multi-party conversations, and recall relevant information from past dialogues. These challenges demand more comprehensive evaluation methods—ones that test a system’s ability to respond within a reasonable time constraint, understand multi-party dialogue contexts, and reason across extended interactions. To address this need, we introduce DialSim, a real-time dialogue simulator designed to evaluate the long-term multi-party dialogue understanding of conversation systems.
@@
DialSim places the conversation system in the role of a main character within a TV show, engaging in extensive conversations based on the show’s scripted content (see Figure 1). During each session, a randomly selected character asks a randomly sampled question at an unpredictable time. The system is evaluated on its ability to respond appropriately, relying solely on the dialogue history and acknowledging when it lacks sufficient information. This approach enables rigorous testing of dialogue comprehension in unpredictable, realistic scenarios. Additionally, the system’s real-time interaction capabilities are assessed through time constraints for responses (e.g., 3s, 5s). Note that DialSim is designed not to evaluate a single LLM, but to evaluate a conversation system’s dialogue understanding ability in an environment as realistic as possible.
@@
To implement DialSim, a dialogue script and corresponding question-answer (QA) pairs are required. For this purpose, we created LongDialQA, a new QA dataset derived from long-term multi-party dialogues. It comprises dialogues from popular TV shows (i.e., Friends, The Big Bang Theory, and The Office), covering 1,300 sessions over five years, totaling around 350,000 tokens. Each session includes more than 1,000 questions curated through two approaches: refining questions from a fan quiz website and generating complex questions using extracted temporal knowledge graphs.
@@
To rigorously evaluate the system’s ability to use dialogue history rather than relying on their pre-trained knowledge, we also constructed adversarial versions of LongDialQA. Since LLM-based conversation systems may possess prior knowledge of the TV shows (see Appendix A), we designed adversarial modifications that alter character names in two ways: by swapping their names with each other (e.g., Joey $\leftrightarrow$ Monica) or by assigning new names to them (e.g., Joey $\rightarrow$ John). These modifications ensure that the system’s responses are grounded only in the contextual dialogue history rather than their pre-trained knowledge.
Our contributions are summarized as follows:
  • We propose DialSim, a novel framework that evaluates the conversation system’s ability to understand long-term, multi-party dialogues under time constraints and unpredictable questioning.
  • DialSim bridges the gap between research and real-world applications by incorporating practical deployment factors (e.g., network latency) in the evaluation of conversation systems, making it a more practical alternative to existing methods that evaluate LLMs rather than systems.
  • We introduce LongDialQA, a new dialogue QA dataset featuring long-term multi-party dialogues with high-quality QA pairs.
  • To ensure robust evaluation and address potential data leakage, we constructed adversarial versions of LongDialQA, guaranteeing that conversation systems are assessed in unseen dialogues.
@@ 2 Related Works
Conversation Systems Evaluation   Early evaluation methods for conversation systems often relied on reference-based metrics (e.g., BLEU Papineni et al. (2002), ROUGE Lin (2004), METEOR Banerjee and Lavie (2005)), which compare model outputs to gold dialogue references but often show weak correlation with human judgment Liu et al. (2016). In contrast, human evaluation—where human annotators assess coherence, factual correctness, consistency, and engagingness of the generated responses—provides reliable assessments Adiwardana et al. (2020); Zhang et al. (2020); Roller et al. (2021); Shuster et al. (2022); Lee et al. (2023), but it is costly and time-consuming.
@@
With the advent of LLMs, new evaluation approaches have emerged. These include having LLMs evaluate utterances directly Li et al. (2023); Kim et al. (2024) or employing platforms (e.g., Chatbot Arena Chiang et al. (2024)) where humans rank responses from different systems. Despite these advances, existing methods are still limited to qualitative assessments of utterances and fail to capture real-world conversational scenarios (e.g., real-time interaction, and long-term multi-party dialogue). To address these limitations, we propose a dialogue simulator, DialSim, designed to evaluate a conversation system’s comprehensive dialogue understanding capabilities.
@@
Datasets Based on TV Show Scripts   While both TV show scripts and other dialogue datasets effectively capture dialogue characteristics, scripts offer a significant advantage due to their abundance and accessibility. This makes them particularly valuable for various dialogue understanding tasks such as question answering (Yang and Choi, 2019; Sang et al., 2022), coreference resolution (Chen et al., 2017; Zhou and Choi, 2018), relation extraction (Rashid and Blanco, 2018; Yu et al., 2020), and summarization (Papalampidi et al., 2020; Chen et al., 2022). While many studies have utilized TV show scripts to create such datasets, only LongDialQA includes unanswerable questions and fully utilizes the extended context.
Dialogue QA Datasets   Dialogue QA datasets are constructed from various sources, capturing diverse conversational settings. Among multi-party dialogue-based QA datasets, Molweni Li et al. (2020) was built on the Ubuntu Chat Corpus Lowe et al. (2015), while MeetingQA Prasad et al. (2023) focuses on meeting dialogues. Additionally, there are script-based dialogue QA datasets, like FriendsQA (Yang and Choi, 2019) and TVShowGuess (Sang et al., 2022), similar to LongDialQA. More recently, LoCoMo (Maharana et al., 2024) introduced a dataset designed to assess an agent’s dialogue comprehension through tasks like event summarization. This dataset was constructed using a combination of LLMs and crowdsourcing. While many datasets have focused on dialogue QA, they either lack multi-party interactions or do not encompass dialogues as long as LongDialQA, which spans 350k tokens. LongDialQA fills this gap by providing an extensive multi-party dialogue corpus for QA tasks.
@@ 3.1.1 Script Preprocessing
The script we used includes 5 consecutive seasons per TV show, with each season containing approximately 20 episodes. Each episode is composed of multiple scenes (i.e., session). Each script includes not only utterances but also descriptions of characters’ actions and scenes, as well as metadata unrelated to the plot (e.g., names of writers and directors). We manually filtered out all irrelevant parts to create $Script_{pre}$, which contains only the conversations between characters. Additionally, since some of our questions involve time conditions (e.g., “Which friend wasn’t allowed to drive Monica’s Porsche in October 1994?”), we manually assigned a date to each scene in $Script_{pre}$ to provide time information to the system. These dates were determined based on the contents of the conversations and the air dates of the episodes. The specific rules for date assignments are detailed in Appendix B. We then selected scenes involving the main character (i.e., Friends: Ross, The Big Bang Theory: Sheldon, The Office: Michael [^2] The characters with the most lines in each script were selected.) from $Script_{pre}$ and sequentially numbered them as sessions $S_{i}$. This process resulted in the final dialogue $\mathcal{D}=\{\mathcal{S}_{1},\mathcal{S}_{2},...,\mathcal{S}_{N}\}$.
Figure 2: The overall process of question generation based on fan quizzes. First, we crawled fan quizzes from the web (1). Next, we applied filtering and revision processes to the crawled data (2). After that, we identified evidence scenes that could provide answers to the questions (3). From this, we created secondary versions of the questions by adding dates to each. We then mapped each question to the scenes by determining whether it is answerable in that scene or not (4). Finally, we applied character style transfer to make the questions more natural (5).
@@ 3.1.2 Fan Quiz-Based Question Generation
We utilized a fan quiz website FunTrivia [^3] https://www.funtrivia.com/ to generate our questions. Fan quizzes cover a range of difficulty levels and focus on major events from each episode, making them promising for evaluating dialogue comprehension. Figure 2 illustrates our process for generating questions using fan quizzes. We began by extracting episode-specific quizzes from the site. Since these quizzes were created by dedicated fans, many required knowledge unrelated to the dialogue itself (e.g., “What is the name of the actor who played the clerk?”). To filter out these questions, we first selected quizzes that could be answered by referencing $Script_{pre}$ using GPT-4 (OpenAI, 2023a). [^4] Fan quizzes exist for each episode, so we annotated them based on $Script_{pre}$ and then matched them to the sessions of $\mathcal{D}$. Questions about scenes without the main character are unanswerable, enabling us to design rigorous tests. Additionally, GPT-4 annotated the scenes that served as evidence for each question. The authors verified these annotations to ensure accurate filtering and scene-mapping.
@@
Figure 3: The overall process of question generation based on the temporal knowledge graph. We first extracted quadruples and constructed a temporal knowledge graph (1). Then, we generated questions based on this and mapped each question to the sessions by determining whether it was answerable in that session or not, similar to fan quiz-based questions (2). Character style transfer was performed afterwards (3).
@@ 3.1.3 Temporal Knowledge Graph-Based Question Generation
Fan quizzes are useful for generating our questions, but since they are episode-specific and user-generated, the questions don’t span multiple episodes and their numbers are limited ($\sim$1K). To address this, we constructed a knowledge graph for each session and used it to generate questions. Initially, we used GPT-4 to extract triples (i.e., [head, relation, tail]) from each session $S_{i}$ in $\mathcal{D}$. These triples were then refined by the authors. We employed 32 relations (e.g., girlfriend) derived from DialogRE (Yu et al., 2020), a high-quality dataset where human annotators manually extracted relations from Friends scripts, classifying relationships between characters into 37 categories. We adapted and modified these relations for our purpose. More details about the relations are provided in Appendix D.1. Finally, we combined the triples from each session with their respective dates to create a temporal knowledge graph (TKG) composed of quadruples (i.e., [head, relation, tail, date]).
@@ 3.1.4 Final Data Processing
Answer Choices Generation    To create multiple-choice questions, we carefully crafted a set of answer choices for each question. First, for all questions, we included a choice “(E) I don’t know.”, which systems must choose if the questions are unanswerable. For questions sourced from fan quizzes, the four answer choices were taken from the original quiz. The correct answers for these questions were the same as the original quiz, while the unanswerable questions were fixed to (E).
@@
Question Style Transfer    In LongDialQA, questions are rephrased to reflect each character’s unique tone, creating the impression that the characters themselves are asking the questions (e.g., Generic style: “How did Rachel buy her new boots?$\rightarrow$ Style of Joey Tribbiani from Friends: “Hey, how did Rachel manage to snag those killer boots, huh?”). This transformation is powered by GPT-4, and subsamples are reviewed by the authors to ensure that the original intent was preserved. More examples of style-transferred questions for each character are in Appendix E.
Adversarial Data Generation   We created two adversarial versions of LongDialQA by altering the names of the characters in two ways: by swapping their names with each other (e.g., Joey $\leftrightarrow$ Monica) or by assigning new names to them (e.g., Joey $\rightarrow$ John). By making these changes, we ensure that all events appearing in dialogues and QA remains consistent, resulting in a fully coherent new dataset.